the length of a rectangular field is 50 feet more than its width. if the perimeter of the basketball court is 300 feet, what is the length of the field
Answer by macston(5194)

W=width; L=length=W+50ft; P=perimeter=300 feet
P=2(L+W)Substitute for L.
300 feet=2(W+50ft+W) Divide each side by 2.
150 feet=2W+50ft Subtract 50 feet from each side.
100 feet=2W Divide each side by 2
50 feet=W The width is 50 feet.
L=W+50ft=50ft+50ft=100 feet ANSWER: The length is 100 feet.
CHECK:
P=2(L+W)
300ft=2(100ft+50ft)
300ft=2(150ft)
300ft=300ft